# Shallow brambles

### v1

@@ 1. Introduction
The equivalence between (1) and (3) is due to Dvořák and Norin [2], while that of (1) and (2) has been shown in [10, Chapter 6]. Having balanced separators of size strongly sublinear in the number of vertices is a very useful property in structural graph theory and algorithm design, which motivates the study of graph classes with polynomial expansion. However, these classes are far from being well understood. Indeed, not much else is known besides the equivalences stated above.
@@
We conclude this introduction by mentioning that the topic of this paper bears some superficial similarities with the Coarse Graph Theory project initiated recently by Georgakopoulos and Papasoglu [5], which aims at revisiting structural graph theory (and in particular the theory of graph minors) from the point of view of metric theory. A fat $K_{t}$-minor or metric $K_{t}$-minor in a graph $G$ in this context is a minor-model of $K_{t}$ in $G$ such that the branch vertices are far apart, and are connected by paths that are also far apart. Here all distances are considered in the graph $G$. This is somewhat orthogonal to the approach we take here, where minor-models have small radius (where distances are computed in the graph induced by the branch vertices, not in the original graph). Nevertheless, this paper can be seen as an effort to understand what remains true if we introduce appropriate distance constraints in classical objects from graph minor theory such as minor-models, brambles, tangles, and well-linked sets.
@@ 4. The proofs
It sufices to show that any $r$-well-linked set $S$ is $(3r,k_{S})$-linked, where $k_{S}:=\lfloor\tfrac{1}{2}\sqrt{|S|}\rfloor$.

### v2

@@ 1. Introduction
The equivalence between (1) and (3) is due to Dvořák and Norin [3], while that of (1) and (2) has been shown in [11, Chapter 6]. Having balanced separators of size strongly sublinear in the number of vertices is a very useful property in structural graph theory and algorithm design, which motivates the study of graph classes with polynomial expansion. However, these classes are far from being well understood. Indeed, not much else is known besides the equivalences stated above.
@@
Related work
The topic of this paper bears some superficial similarities with the Coarse Graph Theory project initiated recently by Georgakopoulos and Papasoglu [6], which aims at revisiting structural graph theory (and in particular the theory of graph minors) from the point of view of metric theory. A fat $K_{t}$-minor or metric $K_{t}$-minor in a graph $G$ in this context is a minor-model of $K_{t}$ in $G$ such that the branch vertices are far apart, and are connected by paths that are also far apart. Here all distances are considered in the graph $G$. A recent paper by Dragan [1] on related topics explores the minimum radius $r$ such that for every bramble $\mathcal{B}$ in a graph $G$, there is a ball of radius $r$ in $G$ that intersects all elements of $\mathcal{B}$. One of the results of [1] is that this minimum radius is within a small multiplicative constant of the tree-length of $G$, the minimum over all tree-decompositions of $G$ of the maximum diameter of a bag.
This is somewhat orthogonal to the approach we take here, where minor-models and bramble elements have small radius (where distances are computed in the graph induced by the branch vertices or elements, not in the original graph). Nevertheless, this paper can be seen as an effort to understand what remains true if we introduce appropriate distance constraints in classical objects from graph minor theory such as minor-models, brambles, tangles, and well-linked sets.
It was pointed out to us by Michał Pilipczuk that our definition of a shallow bramble is related to the cop-width of a graph, a parameter recently introduced by Toruńczyk [18]. Note that very much like brambles (or heavens) in the helicopter version of the cops and robber game of Seymour and Thomas [16], the existence of a radius-$r$ bramble of order $k$ provides a winning strategy for the robber against $k-1$ cops in the version of the game by Toruńczyk when the robber has speed $12r+$ 1.
@@ 4. The proofs
It suffices to show that any $r$-well-linked set $S$ is $(3r,k_{S})$-linked, where $k_{S}:=\lfloor\tfrac{1}{2}\sqrt{|S|}\rfloor$.